# On circumcentered direct methods for monotone variational inequality problems

### v1

@@ 3.2 Convergence Analysis
$\begin{split}\bar{\rho}\sum_{k=k^{\prime}}^{q}\beta_{k}&\leq\sum_{k=k^{\prime}}^{q}\left(\left\lVert x^{k}-\bar{x}\right\rVert^{2}-\left\lVert x^{k+1}-\bar{x}\right\rVert^{2}\right)\\ &\leq\left\lVert x^{k^{\prime}}-\bar{x}\right\rVert^{2}-\left\lVert x^{q+1}-\bar{x}\right\rVert^{2}\\ &\leq\left\lVert x^{k^{\prime}}-\bar{x}\right\rVert^{2}.\end{split}$ (3.31)
Taking limits in (3.31) with $q\to\infty$, we contradict assumption (3.4). Thus, there exists a cluster point of $\{x^{k}\}$ belonging to SOL$(F,C)$.
Finally, we prove that all cluster points of $\{x^{k}\}$ belong to SOL$(F,C)$. Suppose that $\{x^{k}\}$ has a cluster point $z\notin{\rm SOL}(F,C)$. Since SOL$(F,C)$ is closed by Proposition 2.7, and $\lim_{k\to\infty}\left\lVert x^{k+1}-x^{k}\right\rVert=0$ by Lemma 3.9(iii), we invoke Proposition 2.10 to obtain a subsequence $\{x^{j_{k}}\}$ of $\{x^{k}\}$ and a real number $\sigma>0$ such that
${\rm dist}\left(x^{j_{k}},{\rm SOL}(F,C)\right)>\sigma,$ (3.32)
@@
${\rm dist}\left(x^{j_{k}+1},{\rm SOL}(F,C)\right)>{\rm dist}\left(x^{j_{k}},{\rm SOL}(F,C)\right).$ (3.33)
Take $\gamma_{k}$ as defined in (3.26). Note that for $\bar{x}\in{\rm SOL}(F,C)$, $\{\gamma_{k}(\bar{x})\}$ is bounded by Lemma 3.9(ii). Define $\gamma:{\rm SOL}(F,C)\to\mathds{R}$ as
$\gamma(x):=\liminf_{k\rightarrow\infty}\gamma_{j_{k}}(x).$ (3.34)
We claim that $\gamma(x)>0$ for all $x\in{\rm SOL}(F,C)$. Otherwise, By Proposition 2.8 $\{x^{j_{k}}\}$ has a cluster point in SOL$(F,C)$, in contradiction with (3.32). We claim now that $\gamma$ is continuous in SOL(F,C). Take $x,x^{\prime}\in{\rm SOL}(F,C)$. Note that
@@
Let $V$ be the set of cluster points of $\{x^{k}\}$. We have shown above that $V\cap{\rm SOL(F,C)}\neq\emptyset$. Since $\{x^{k}\}$ is bounded, $V$ is compact and so is $V\cap{\rm SOL}(F,C)$. It follows that $\gamma$ attains its minimum on $V\cap{\rm SOL}(F,C)$ at some $x^{*}$, so that $\gamma(x)\geq\gamma(x^{*})>0$ for all $x\in V\cap{\rm SOL}(F,C)$, using the claim above.
@@
$\gamma_{j_{k}}(x)\geq\frac{\gamma(x^{*})}{2},$ (3.35)
@@
$\beta_{j_{k}}<\frac{\gamma(x^{*})}{\theta},$ (3.36)
for all $k\geq\hat{k}$. Note that $\hat{k}$ exists because, for large enough $k$, (3.35) holds by virtue of (3.34), and (3.36) holds because $\lim_{k\to\infty}\beta_{k}=0$. In view of (3.27), we get, for all $x\in V\cap{\rm SOL}(F,C)$ and all $k\geq\hat{k}$,
$\begin{split}\left\lVert x^{j_{k}+1}-x\right\rVert^{2}&\leq\left\lVert x^{j_{k}}-x\right\rVert^{2}-\beta_{j_{k}}\left(2\frac{\gamma_{j_{k}}(x)}{\eta_{j_{k}}}-\beta_{j_{k}}\right)\\ &\leq\left\lVert x^{j_{k}}-x\right\rVert^{2}-\beta_{j_{k}}\left(2\frac{\gamma_{j_{k}}(x)}{\theta}-\beta_{j_{k}}\right)\\ &<\left\lVert x^{j_{k}}-x\right\rVert^{2},\end{split}$ (3.37)
@@
${\rm dist}(x^{j_{k}+1},V\cap{\rm SOL}(F,C))\leq{\rm dist}(x^{j_{k}},V\cap{\rm SOL}(F,C)),$
for all $k\geq\hat{k}$, in contradiction with 3.33. The contradiction arises from assuming that $\{x^{k}\}$ has cluster points out of SOL$(F,C)$, and therefore all cluster points of $\{x^{k}\}$ solve VI$(F,C)$. ∎
We summarize the convergence sequence properties of Algorithm 1 in the following corollary.
@@
If $T$ is paramonotone and coercive (in the sense of Assumption 3.5), and ${\rm SOL(F,C)}\neq\emptyset$, then any sequence $\{x^{k}\}$ generated by Algorithm 1 is bounded, $\lim_{k\to\infty}\left\lVert x^{k+1}-x^{k}\right\rVert=0$, and all cluster points of $\{x^{k}\}$ belong to SOL$(F,C)$. If VIP$(F,C)$ has a unique solution then the whole sequence $\{x^{k}\}$ converges to it.
@@
3.3 Algorithm 1 when the $C_{i}$’s are sublevel sets of convex functions
@@
$C_{i}=\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\},$ (3.38)
where each $g_{i}:\mathds{R}^{n}\to\mathds{R}$ is convex. In principle, this form entails no loss of generality, because we can take $g_{i}(x)=\left\lVert P_{C_{i}}(x)-x\right\rVert^{2}$, but this case turns out to be computationally advantageous only when we have easily computable formulae for $g_{i}$ and its subdifferential. We remark that this is indeed the case in a very large number of applications. When the $C_{i}$’s are as in (3.38), we can use a very convenient operator $S^{i}$, which leads to explicit closed formulae for the iteration of Algorithm 1.
We define a separator operator for the set $C_{i}$ as
$S^{i}(x)=\{y\in\mathds{R}^{n}:g_{i}(x)+\langle u^{i},y-x\rangle\leq 0\},$ (3.39)
where $u^{i}$ is a subgradient of $g_{i}$ at $x$.
In connection with the sequence $\{x^{k}\}$ generated by Algorithm 1 we define $S^{i}_{k}:=S_{i}(x^{k})$, i.e.,
$S^{i}_{k}=\{y\in\mathds{R}^{n}:g_{i}(x^{k})+\langle u^{ik},y-x^{k}\rangle\leq 0\},$ (3.40)
where now $u^{ik}$ is a subgradient of $g_{i}$ at $x^{k}$. Finally, we define ${\mathbf{S}}_{k}\subset\mathds{R}^{nm}$ as
${\mathbf{S}}_{k}=S^{1}_{k}\times\dots\times S^{m}_{k}.$ (3.41)
This choice of the set ${\mathbf{S}}_{k}$ leads to the closed formulae presentation of Algorithm 1, namely:
$x^{k+1}=T_{\textbf{S}_{k},\textbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)=x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})-\alpha_{k}w^{k},$ (3.42)
@@
$\alpha_{k}=\frac{\sum_{i=1}^{m}\left\lVert v^{ik}\right\rVert^{2}}{m\left\lVert w^{k}\right\rVert^{2}},$ (3.43)
$w^{k}=\frac{1}{m}\sum_{i=1}^{m}v^{ik},$ (3.44)
$v^{ik}=\left[\frac{\max\left\{0,g_{i}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)\right\}}{\left\lVert u^{ik}\right\rVert}\right]u^{ik},$ (3.45)
with $u^{ik}\in\partial g_{i}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)$.
@@
We present some comments on the formulae (3.42)–(3.45).
  • i)
    The formula of $\alpha_{k}$ follows from the definition of the circumcenter, using Proposition 2.2 in (3.45); the computation of $\alpha_{k}$ is similar to the one presented in [11], in the definition of the PACA Algorithm for solving a CFP. In fact, the main difference between our closed formulae above and those in [11] is the presence of the term $(\beta/\eta_{k})F(x^{k})$ in (3.42), related to the operator $F$, which is specific of VIP$(F,C)$ and is absent in CFP.
  • ii)
    The closed formulae (3.42)-(3.45) were not used in the convergence analysis of Subsection 3.2, which is performed for general separating sets satisfying the conditions in Definition 1.2, but we note that they are essential for the effective implementation of the algorithm.
  • iii)
@@
  • iv)
    Observe that $\partial g_{i}(x)\neq\emptyset$ for all $x\in\mathds{R}^{n}$, because we assume that $g_{i}$ is convex and dom$(g_{i})=\mathds{R}^{n}$ for all $i$.
In order to ensure that the convergence analysis in Subsection 3.2 holds for ${\mathbf{S}}_{k}$ as defined by (3.40), (3.41), we must prove that the separating operator $S^{i}$ defined by (3.39) satisfies indeed the conditions in Definition 1.2. We establish this in the next two propositions.
@@
Let $f:\mathds{R}^{n}\to\mathds{R}$ be a convex function and $V\subset\mathds{R}^{n}$a nonempty, convex and compact set. Define $C=\{z\in\mathds{R}^{n}:f(z)\leq 0\}$. Given $x\in\mathds{R}^{n}$ and $v\in\partial f(x)$, define $\bar{S}(x,u):=\{z\in\mathds{R}^{n}|f(x)+\langle u,z-x\rangle\leq 0\}$. Then, there exists $\kappa>0$ such that
${\rm dist}(P_{\bar{S}(x,v)\cap X}(x),C)\leq\kappa\,{\rm dist}(x,C).$ (3.46)
@@
Given a convex function $g:\mathds{R}^{n}\to\mathds{R}\,\,(1\leq i\leq m)$ define $C:=\{x\in\mathds{R}^{n}:g(x)\leq 0\}$. Define $S(x)\subset\mathds{R}^{n}$ as
$S(x):=\{y\in\mathds{R}^{n}:g(x)+\langle u,y-x\rangle\leq 0\},$
where $u\in\mathds{R}^{n}$ is a subgradient of $g$ at $x$. Then $S$ is a separating set for $C$ in the sense of Definition 1.2.
@@
The fact that $S$ satisfies conditions (i)-(iii) in Definition 1.2 has been proved in Proposition 2.8 of [3]. We now deal with condition (iv). The result is an immediate consequence of Proposition 3.13. ∎
Proposition 3.14 establishes that ${\mathbf{S}}_{k}$, as defined in (3.41), fits the convergence analysis of Algorithm 1 carried out in Subsection 3.2.
We mention that an argument similar to the one in the proof of Proposition 3.14 establishes that the operator ${\mathbf{S}}:\mathds{R}^{nm}\to{\mathcal{P}}(\mathds{R}^{nm})$ given by ${\mathbf{S}}({\mathbf{x}})=(S^{1}(x^{1}),\dots,S^{m}(x^{m}))$, where ${\mathbf{x}}=(x^{1},\dots,x^{m})$ with $x^{i}\in\mathds{R}^{n}\,\,(1\leq i\leq m)$, and $S^{i}$ is defined as in (3.39), satisfies conditions (i)-(iv) in Definition 1.2. The proof for conditions (i)-(iii) can be found in [3]. However, this result is not needed for our convergence analysis.
4 An explicit circumcenter method for monotone variational inequality problem
In this section, we introduce an algorithm that utilizes approximate projections (projections onto halfspaces) instead of the exact projection. Different from the algorithm in Section 3, this algorithm only requires the assumption of monotonicity instead of paramonotonicity. First, we introduce the Explicit Circumcenter Method for monotone variational inequalities.
Algorithm 2 Explicit Circumcenter Method (ECM) for ${\rm VIP}(F,C)$ with monotone operators
Initialization:
Fix an exogenous constant $\theta>0$ and define
$x^{0}:=0~{}and~{}z^{0}\in\mathds{R}^{n}$
Choose a stepsize sequence $\{\beta_{k}\}$ satisfying:
$\sum_{k=0}^{\infty}\beta_{k}=\infty,$ (4.1)
and
$\sum_{k=0}^{\infty}\beta_{k}^{2}<\infty,$ (4.2)
Iterative Step:
Given $z^{k}$, if $g_{i}(z^{k})\leq 0$ for all $i\in\{1,\ldots,m\}$, then take $\tilde{y}^{k}:=z^{k}$. Else, perform the following inner loop, generating points $y^{k,0}$, $y^{k,1}$, $\ldots$. Take $y^{k,0}=z^{k}$, define
$y^{k,j+1}:=T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j}).$ (4.3)
Stop the inner loop when $j=j(k)$ , defined as
$j(k):=\min\left\{j\geq 0:\frac{g(y^{k,j})\|y^{k,j}-w\|}{g(y^{k,j})-g(w)}\leq\theta\beta_{k}\right\},$ (4.4)

### v2

@@ 3.2. Convergence Analysis
(3.27) $\begin{split}\bar{\rho}\sum_{k=k^{\prime}}^{q}\beta_{k}&\leq\sum_{k=k^{\prime}}^{q}\left(\left\lVert x^{k}-\bar{x}\right\rVert^{2}-\left\lVert x^{k+1}-\bar{x}\right\rVert^{2}\right)\\ &\leq\left\lVert x^{k^{\prime}}-\bar{x}\right\rVert^{2}-\left\lVert x^{q+1}-\bar{x}\right\rVert^{2}\\ &\leq\left\lVert x^{k^{\prime}}-\bar{x}\right\rVert^{2}.\end{split}$
Taking limits in (3.27) with $q\to\infty$, we contradict assumption (1.22). Thus, there exists a cluster point of $\{x^{k}\}$ belonging to $\operatorname{SOL}(F,C)$.
Finally, we prove that all cluster points of $\{x^{k}\}$ belong to $\operatorname{SOL}(F,C)$. Suppose that $\{x^{k}\}$ has a cluster point $z\notin\operatorname{SOL}(F,C)$. Since $\operatorname{SOL}(F,C)$ is closed by Proposition 2.7, and $\lim_{k\to\infty}\left\lVert x^{k+1}-x^{k}\right\rVert=0$ by Lemma 3.4(iii), we invoke Proposition 2.10 to obtain a subsequence $\{x^{j_{k}}\}$ of $\{x^{k}\}$ and a real number $\sigma>0$ such that
(3.28) $\operatorname*{dist}\left(x^{j_{k}},\operatorname{SOL}(F,C)\right)>\sigma,$
@@
(3.29) $\operatorname*{dist}\left(x^{j_{k}+1},\operatorname{SOL}(F,C)\right)>\operatorname*{dist}\left(x^{j_{k}},\operatorname{SOL}(F,C)\right).$
Take $\gamma_{k}$ as defined in (3.22). Note that for any $\bar{x}\in\operatorname{SOL}(F,C)$, $\{\gamma_{k}(\bar{x})\}$ is bounded by Lemma 3.4(ii). Define $\gamma:\operatorname{SOL}(F,C)\to\mathds{R}$ as
(3.30) $\gamma(x)\coloneqq\liminf_{k\rightarrow\infty}\gamma_{j_{k}}(x).$
We claim that $\gamma(x)>0$ for all $x\in\operatorname{SOL}(F,C)$. Otherwise, by Proposition 2.8 $\{x^{j_{k}}\}$ has a cluster point in $\operatorname{SOL}(F,C)$, in contradiction with (3.28). We next claim that $\gamma$ is continuous on $\operatorname{SOL}(F,C)$. Take $x,x^{\prime}\in\operatorname{SOL}(F,C)$. Note that
@@
Let $V$ be the set of cluster points of $\{x^{k}\}$. We have shown above that $V\cap{\operatorname{SOL}(F,C)}\neq\emptyset$. Since $\{x^{k}\}$ is bounded, $V$ is compact and so is $V\cap\operatorname{SOL}(F,C)$. It follows that $\gamma$ attains its minimum on $V\cap\operatorname{SOL}(F,C)$ at some $x^{*}$, so that $\gamma(x)\geq\gamma(x^{*})>0$ for all $x\in V\cap\operatorname{SOL}(F,C)$, using the claim above.
@@
(3.31) $\gamma_{j_{k}}(x)\geq\frac{\gamma(x^{*})}{2},$
@@
(3.32) $\beta_{j_{k}}<\frac{\gamma(x^{*})}{\theta},$
for all $k\geq\hat{k}$. Note that $\hat{k}$ exists because, for large enough $k$, (3.31) holds by virtue of (3.30), and (3.32) holds because $\lim_{k\to\infty}\beta_{k}=0$. In view of (3.23), we get, for all $x\in V\cap\operatorname{SOL}(F,C)$ and all $k\geq\hat{k}$,
(3.33) $\begin{split}\left\lVert x^{j_{k}+1}-x\right\rVert^{2}&\leq\left\lVert x^{j_{k}}-x\right\rVert^{2}-\beta_{j_{k}}\left(2\frac{\gamma_{j_{k}}(x)}{\eta_{j_{k}}}-\beta_{j_{k}}\right)\\ &\leq\left\lVert x^{j_{k}}-x\right\rVert^{2}-\beta_{j_{k}}\left(2\frac{\gamma_{j_{k}}(x)}{\theta}-\beta_{j_{k}}\right)\\ &<\left\lVert x^{j_{k}}-x\right\rVert^{2},\end{split}$
@@
$\operatorname*{dist}(x^{j_{k}+1},V\cap\operatorname{SOL}(F,C))\leq\operatorname*{dist}(x^{j_{k}},V\cap\operatorname{SOL}(F,C)),$
for all $k\geq\hat{k}$, in contradiction with (3.29). The contradiction arises from assuming that $\{x^{k}\}$ has cluster points outside $\operatorname{SOL}(F,C)$, and therefore all cluster points of $\{x^{k}\}$ solve $\operatorname{VIP}(F,C)$. ∎
We summarize the convergence properties of the sequence generated by Algorithm 1 in the following corollary.
@@
If $F$ and $C$ satisfy Assumptions 1.5, 1.7 and 1.9, (i.e. $\operatorname{SOL}(F,C)$$\neq\emptyset$ and $F$ is continuous and paramonotone), then any sequence $\{x^{k}\}$ generated by Algorithm 1 is bounded, $\lim_{k\to\infty}\left\lVert x^{k+1}-x^{k}\right\rVert=0$, and all cluster points of $\{x^{k}\}$ belong to $\operatorname{SOL}(F,C)$. If $\operatorname{VIP}(F,C)$ has a unique solution, then the whole sequence $\{x^{k}\}$ converges to it.
@@
3.3. Algorithm 1 when the $C_{i}$’s are sublevel sets of convex functions
@@
(3.34) $C_{i}=\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\},$
where each $g_{i}:\mathds{R}^{n}\to\mathds{R}$ is convex. In principle, this form entails no loss of generality, because we can take $g_{i}(x)=\left\lVert P_{C_{i}}(x)-x\right\rVert^{2}$, but this case turns out to be computationally advantageous only when we have easily computable formulae for $g_{i}$ and its subdifferential. We remark that this is indeed the case in a large number of applications. When the $C_{i}$’s are as in (3.34), we can use a very convenient operator $S^{i}$, which leads to explicit closed-form expressions for the iteration of Algorithm 1.
We define a separating operator for the set $C_{i}$ as
(3.35) $S^{i}(x)=\{y\in\mathds{R}^{n}:g_{i}(x)+\langle u_{i},y-x\rangle\leq 0\},$
where $u_{i}$ is a subgradient of $g_{i}$ at $x$.
In connection with the sequence $\{x^{k}\}$ generated by Algorithm 1 we define $S^{i}_{k}\coloneqq S_{i}(x^{k})$, i.e.,
(3.36) $S^{i}_{k}=\{y\in\mathds{R}^{n}:g_{i}(x^{k})+\langle u^{k}_{i},y-x^{k}\rangle\leq 0\},$
where now $u^{k}_{i}$ is a subgradient of $g_{i}$ at $x^{k}$. Finally, we define ${\mathbf{S}}_{k}\subset\mathds{R}^{nm}$ as
(3.37) ${\mathbf{S}}_{k}=S^{1}_{k}\times\dots\times S^{m}_{k}.$
This choice of the set ${\mathbf{S}}_{k}$ leads to the closed-form presentation of Algorithm 1, namely:
(3.38) $x^{k+1}=T_{\textbf{S}_{k},\textbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)=x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})-\alpha_{k}w^{k},$
@@
(3.39) $\alpha_{k}=\frac{\sum_{i=1}^{m}\left\lVert v^{k}_{i}\right\rVert^{2}}{m\left\lVert w^{k}\right\rVert^{2}},$
(3.40) $w^{k}=\frac{1}{m}\sum_{i=1}^{m}v^{k}_{i},$
(3.41) $v^{k}_{i}=\left[\frac{\max\left\{0,g_{i}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)\right\}}{\left\lVert u^{k}_{i}\right\rVert}\right]u^{k}_{i},\text{ for }i=1,2,\dots,m,$
with $u^{k}_{i}\in\partial g_{i}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)$.
@@
We now make the following comments on formulae (3.38), (3.39), (3.40), and (3.41).
  1. (i)
    The formula for $\alpha_{k}$ follows from the definition of the circumcenter, using Proposition 2.2 in (3.41); the computation of $\alpha_{k}$ is similar to the one presented in [10], in the definition of the PACA Algorithm for solving CFP. In fact, the main difference between our closed formulae above and those in [10] is the presence of the term $(\beta/\eta_{k})F(x^{k})$ in (3.38), related to the operator $F$, which is specific to $\operatorname{VIP}(F,C)$ and is absent in CFPs.
  2. (ii)
    The closed formulae (3.38)-(3.41) were not used in the convergence analysis of Subsection 3.2, which deals with general separating operators satisfying the conditions in Definition 1.4, but we note that they are essential for the effective implementation of the algorithm.
  3. (iii)
@@
  4. (iv)
    Observe that $\partial g_{i}(x)\neq\emptyset$ for all $x\in\mathds{R}^{n}$, because we assume that $g_{i}$ is convex and $\operatorname{dom}(g_{i})=\mathds{R}^{n}$ for all $i$.
In order to ensure that the convergence analysis in Subsection 3.2 holds for ${\mathbf{S}}_{k}$ as defined by (3.36) and (3.37), we must prove that the separating operator $S^{i}$ defined by (3.35) indeed satisfies the conditions in Definition 1.4. We establish this in the next two propositions.
@@
Let $f:\mathds{R}^{n}\to\mathds{R}$ be a convex function and $V\subset\mathds{R}^{n}$ be a nonempty, convex, and compact set. Define $C=\{z\in\mathds{R}^{n}:f(z)\leq 0\}$. Given $x\in\mathds{R}^{n}$ and $v\in\partial f(x)$, define $\bar{S}(x,u)\coloneqq\{z\in\mathds{R}^{n}:f(x)+\langle u,z-x\rangle\leq 0\}$. Then, there exists $\kappa\in[0,1)$ such that
(3.42) $\operatorname*{dist}(P_{\bar{S}(x,v)\cap X}(x),C)\leq\kappa\,\operatorname*{dist}(x,C),$
@@
Given a convex function $g:\mathds{R}^{n}\to\mathds{R}$, let $C\coloneqq\{x\in\mathds{R}^{n}:g(x)\leq 0\}$. Define $S(x)\subset\mathds{R}^{n}$ by
$S(x)\coloneqq\{y\in\mathds{R}^{n}:g(x)+\langle u,y-x\rangle\leq 0\},$
where $u\in\mathds{R}^{n}$ is a subgradient of $g$ at $x$. Then $S$ is a separating operator for $C$ in the sense of Definition 1.4.
@@
The fact that $S$ satisfies conditions (i)-(iii) in Definition 1.4 has been proved in Proposition 2.8 of [3]. We now verify condition (iv). The result is an immediate consequence of Proposition 3.8. ∎
Proposition 3.9 establishes that ${\mathbf{S}}_{k}$, as defined in (3.37), is compatible with the convergence analysis of Algorithm 1 carried out in Subsection 3.2.
We also mention that an argument similar to the one in the proof of Proposition 3.9 establishes that the operator ${\mathbf{S}}:\mathds{R}^{nm}\to{\mathcal{P}}(\mathds{R}^{nm})$ given by ${\mathbf{S}}({\mathbf{x}})=(S^{1}(x^{1}),\dots,S^{m}(x^{m}))$, where ${\mathbf{x}}=(x^{1},\dots,x^{m})$ with $x^{i}\in\mathds{R}^{n}\,\,(1\leq i\leq m)$, and $S^{i}$ is defined as in (3.35), satisfies conditions (i)-(iv) in Definition 1.4. The proof for conditions (i)-(iii) can be found in [3]. However, this result is not needed for our convergence analysis.
4. An explicit circumcenter method for the monotone variational inequality problem
In this section, we introduce an algorithm that unlike the algorithm in Section 3, requires only the assumption of monotonicity instead of paramonotonicity. Our proposal is formally described in Algorithm 2 below. It is an acceleration of the method developed by Bello-Cruz and Iusem [18]. The main difference between our proposal is the use of circumcenters.
To the best of our knowledge, there are few algorithms that solve a VIP with a monotone operator using approximate projections that may produce infeasible iterates. An algorithm with inexact projections was proposed in [28], but those projections remain feasible and still require solving an auxiliary problem at each iteration. In contrast, the approximate projections we implement in this paper are computationally cheaper, even though they may yield infeasible iterates.
Algorithm 2 Explicit Circumcenter Method (ECM) for $\operatorname{VIP}(F,C)$ with monotone operators
Slater point $w\in\mathds{R}^{n}$, that is, a point such that $g_{i}(w)<0$ for all $i\in\{1,\ldots,m\}$; exogenous constant $\theta>0$; sequence $\{\beta_{k}\}$ as defined in Assumption 1.6.
Initialization: Set $x^{0}:=0$ and choose $z^{0}\in\mathds{R}^{n}$.
for $k=0,1,2,\dots$ do $\triangleright$ Iterative Step
  Define $g(x)\coloneqq\max_{1\leq i\leq m}g_{i}(x)$.
  if $g_{i}(z^{k})\leq 0$ for all $i\in\{1,\ldots,m\}$ then
   $\tilde{y}^{k}\leftarrow z^{k}$
[…]
   Set $y^{k,0}\leftarrow z^{k}$ and $j\leftarrow 0$.
   repeat
     Compute next inner iterate:
(4.1) $y^{k,j+1}:=T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j}).$
     Check stopping condition for $j$:
(4.2) $\text{Condition: }\frac{g(y^{k,j})\|y^{k,j}-w\|}{g(y^{k,j})-g(w)}\leq\theta\beta_{k},$